- 17 Mar 2013 [10:25]
- ADJUST: +3500 ms/day to 34000 which is a big change required after last night and the brutal hacking (during which the pendulum was stopped and jolted about)
- 17 Mar 2013 [01:45]
- Winders and witnesses - a Troy of Helens: Helen Arney @helenarney, Helen Scales, Helen Czerski @helenczerski, and the non-Helens: Jasmin Chana, Michael Conterio @Sci_Cam, Matt Parker @standupmaths
- 17 Mar 2013 [01:41]
- GLITCH Major changes tonight - opened up the pendulum chamber with a saw (!) to improve air flow. Hope it will work.